OpenStudy (anonymous):

Solve: 6x − 10 + 14 = 22 4 times 1 over 3 6 3 2

OpenStudy (lncognlto):

So which do you think the answer is? :)

OpenStudy (anonymous):

second one

OpenStudy (lncognlto):

6x - 10 + 14 = 22 6x = 22 - 14 + 10 6x = 18 x = 3